# Strategy for Direct Detection of Chiral Phonons with Phase-Structured Free Electrons

### v1

@@ Abstract
Chirality plays a central role in nearly all aspects of physics, chemistry, and biology. At large length scales, chiral gravitational waves have been considered as sources of circularly polarized contributions to the cosmic microwave background [1], while at short length scales chirality dictates molecular interactions and organizations underpinning biological processes. In condensed matter physics, chirality and spin-momentum locking are at the heart of the menagerie of identified Hall effects, as well as stable [2, 3] and transient free space [4, 5] and evanescent [6, 7, 8, 9, 10] topological electromagnetic field textures. Following the report [11] of helicity-resolved Raman scattering experiments from transition-metal dichalcogenide (TMD) atomic layers, it was predicted [12] that two-dimensional (2D) atomic crystals with hexagonal symmetry can host chiral phonons with well-defined pseudo angular momentum (PAM) arising from discrete rotational crystal symmetry. Direct experimental evidence of chiral phonons was subsequently reported based on transient infrared circular dichroism measurements of a monolayer TMD [13], generating significant interest in the potential for leveraging chiral phonons to achieve novel material properties that are sensitive to low-energy spin-selective excitations. Beyond 2D hexagonal lattices, chiral phonons have been identified in other 2D [14, 15, 16, 17, 18, 19] and 3D [20, 21] crystalline systems, as well as in bio-organic molecules [22]. Chiral phonons have also been recently identified as participants in various exotic condensed matter processes including anomalous thermal Hall conductivity [23] and the chiral-phonon-activated spin Seebeck effect [24], among others [25, 26]. Despite the high-level of current interest in chiral phonons, their experimental detection has been largely limited to optical Raman scattering measurements, with spatial resolution and linear momentum limited by the photon wavelength, and direct detection remains an outstanding challenge impeding further progress [27].
@@
Meanwhile, instrumental advances during the past decade in aberration correction, monochromation, and detector technologies have made it possible to perform vibrational spectroscopy inside a scanning transmission electron microscope (STEM) [28, 29, 30, 31, 32]. In particular, momentum-resolved electron energy loss ($q$-EEL) spectroscopy [29, 33, 34, 35], and the related high-resolution EEL spectroscopy (HREELS) technique [36, 37], have been used to measure the energy-momentum dispersion of phonons in mono- and few-layer 2D atomic crystals. While these measurements have demonstrated the ability of inelastic free electron scattering to complement other phonon measurement techniques, such as inelastic neutron and x-ray scattering, the absence of an obvious electron polarization degree of freedom has hindered the spin-selective detection of chiral phonons using free electrons. However, the transverse phase shaping and coherent detection schemes that were originally developed for core-loss EEL spectroscopy and that enabled magnetic circular dichroism measurements in the TEM [38, 39, 40] have been recently adapted to the low loss regime. By selectively monitoring inelastic scattering of free electrons between states with tailored transverse phase profiles, energy loss processes involving desired target mode symmetries [41] and OAM states [42, 43] have been demonstrated, and optical polarization analogs in free electron scattering have been identified [44, 45]. The full potential of these measurement schemes is currently being explored, while the transverse phase shaping and selection technologies underlying them are actively under development [46, 47, 48, 49, 50].
@@
In this Letter we provide a strategy for leveraging inelastic scattering of transverse phase shaped free electrons to achieve direct detection of chiral phonons within a STEM. We introduce pinwheel free electron states composed of a small number of phased plane wave components that share the same reciprocal space three-fold discrete rotational symmetry as chiral phonons in hexagonal monolayer 2D atomic crystals at the K and $\textrm{K}^{\prime}$ points, which introduces a new PAM degree of freedom to the probing electrons. In analogy to previous work on OAM-resolved EEL spectroscopy [51, 42, 52, 43, 44, 53, 45], pre- and post-selection of the transverse free electron states enables tracking of PAM exchange between pinwheel free electrons and the sample. We trace the utility of the identified phase-shaped free electron transitions to the topological spin textures of their associated transition electric fields, which can be tailored to selectively couple to chiral phonons of the same symmetry. Compared to light-based probes, inelastic electron scattering observables offer superior spatial resolution, in addition to an often simpler interpretation since the Brillouin zone (BZ) edges can be accessed directly. This work informs current and future efforts to detect and characterize chiral phonons using inelastic free electron scattering, and identifies specific experimental geometries and parameters to achieve detection.
@@
Figure 1: hBN phonons and $q$-EEL spectroscopy. a) Scheme of $q$-EEL process involving momentum transfer between initial $\mathbf{k}_{i}$ (gray) and final $\mathbf{k}_{f}$ (red) free electron states. The recoil momentum $\hbar{\bf q}$ has projection ${\bf q}_{\parallel}=({\bf q}\cdot\hat{\bf x})\hat{\bf x}$ onto the 2D sample situated in the $xy$ plane. b) Calculated hBN phonon eigenmode dispersion (gray) along $^{\prime}\to\Gamma\to\textrm{K}$ K and normalized DDCS (color map). c) Real-space in-plane phonon eigenmodes at K and $\textrm{K}^{\prime}$. Pink circles indicate circular motion of sublattices, while arrows show the direction of rotation. Integer PAM is noted in the lower-right of each panel. d) Lineouts of the DDCS spectra at K and $\textrm{K}^{\prime}$ from panel (b). All calculations include an empirical damping rate of $\eta=5$ meV/$\textrm{meV}/\hbar$ at zero temperature and a vacuum background.
@@
As a model 2D atomic crystal with hexagonal symmetry, we consider chiral phonons in monolayer hexagonal boron nitride (hBN) situated in the $xy$ plane as dipticted in Fig. 1a. Eigenenergies $\hbar\omega_{\lambda}(\mathbf{k}_{\parallel})$ and eigenvectors $\bm{\xi}_{\lambda}(\mathbf{k}_{\parallel})$ of the six hBN phonon bands, indexed by $\lambda$, are computed within the harmonic crystal approximation [54] with force constants taken from the literature [55, 56, 57]. The gray lines in Fig. 1b show the eigenenergies as a function of $\mathbf{k}_{\parallel}$ within the first BZ along $^{\prime}\to\Gamma\to\textrm{K}$ K. In contrast to graphene, the lack of inversion symmetry within the 2D plane and the fact that time reversal symmetry is broken at the K and $\textrm{K}^{\prime}$ points leads to non-degenerate in-plane valley phonon modes at the K and $\textrm{K}^{\prime}$ reciprocal space points with well-defined PAM [12]. Real-space motion of the four in-plane phonon eigenmodes at the K and $\textrm{K}^{\prime}$ points are shown in Fig. 1c. The lowest and highest energy in-plane modes involve opposing circular motion of both sublattices, with total PAM $=0\hbar$. For the other two in-plane modes at $119$ meV and $135$ meV, while one sublattice remains motionless, the other executes circular motion, the handedness of which determines whether the PAM =±$=\pm 1\hbar$. The depicted phonons exhibit spin-momentum locking in that the $z$-oriented spin associated with each sublattice (equivalent to handedness of circular motion) changes sign under linear momentum reversal, i.e., $\textrm{K}\leftrightarrow\textrm{K}^{\prime}$.
@@
For an incoming probing electron with speed $v_{i}$, the state-resolved loss rate describing inelastic scattering at first order from the 2D atomic crystal at loss energy $\hbar\omega_{if}$ is [58]
@@
where $L^{3}$ defines the probing electron’s quantization volume, $\Omega_{0}$ is the unit cell volume, and $\mathbf{k}$ is within the first BZ. The quantity ~$\tilde{\mathcal{F}}_{\lambda}(\mathbf{k})=\sum_{\mathbf{G}}\sum_{\kappa}\bm{\xi}^{*}_{\lambda\kappa}\cdot M_{\kappa}^{-1/2}\cdot e\bar{\mathbf{Z}}_{\kappa}(\mathbf{k}+\mathbf{G})\cdot\mathbf{E}_{fi,\mathbf{k}+\mathbf{G}\kappa}^{0}$ involves the atomic masses $M_{\kappa}$ and effective charges $eZ_{\kappa}$ of each sublattice, indexed by $\kappa$, and $\mathbf{G}$ denotes a reciprocal lattice vector. $\mathbf{E}_{fi,\mathbf{k}+\mathbf{G}\kappa}^{0}$ is the Fourier coefficient of the quasistatic transition electric field $\mathbf{E}^{0}_{fi}(\mathbf{k})=-i\mathbf{k}(4\pi/k^{2})\,\rho_{fi}(\mathbf{k})$, where $\rho_{fi}(\mathbf{k})$ is the Fourier transform of the charge density $\rho_{fi}(\mathbf{x})=-e\psi^{*}_{f}(\mathbf{x})\psi_{i}(\mathbf{x})$ associated with the free electron transition from initial state $\psi_{i}(\mathbf{x})$ to final state $\psi_{f}(\mathbf{x})$. A phenomenological damping parameter $\eta$ is introduced to give spectral features finite width. Fig. 1a shows a generic wide field inelastic electron scattering event as typically measured by $q$-EELS whereby an incoming free electron plane wave with wave vector $\mathbf{k}_{i}$ (gray) scatters to an outgoing plane wave with wave vector $\mathbf{k}_{f}$ (red) via interaction with the atomic crystal. In the low loss energy range considered, the linear momentum $\hbar\mathbf{q}=\hbar(\mathbf{k}_{i}-\mathbf{k}_{f})$ lost by the probing electron is transferred into excited phonons with Bloch vector $\mathbf{k}_{\parallel}$. The transition field associated with this situation is
@@
which is linearly polarized along the recoil direction $\hat{\mathbf{q}}$. The rate of phonon mode excitation is dictated by ~$\tilde{\mathcal{F}}_{\lambda}(\mathbf{q})\propto\bm{\xi}^{*}_{\lambda\kappa}\cdot\mathbf{E}^{0}_{\mathbf{q},\mathbf{k}+\mathbf{G}\kappa}\propto\bm{\xi}^{*}_{\lambda\kappa}\cdot\hat{\mathbf{q}}$. As a consequence, the linearly polarized transition field couples identically to circularly polarized in-plane phonons, and it is not possible to distinguish the PAM of the phonon modes at K and $\textrm{K}^{\prime}$ points from $q$-EEL spectra alone. Indeed, spectral lineouts of the double differential cross section (DDCS) [59, 60] at the K and $\textrm{K}^{\prime}$ points from Fig. 1b are shown in Fig. 1d to be indistinguishable.
Figure 2: Transition electric fields involving three-fold pinwheel states. a) Scheme showing a transition from the $|P=3,k_{\parallel},\sigma=+1,\phi_{0}\rangle$ pinwheel state to $|\mathbf{k}^{\parallel}_{f}=\mathbf{0}\rangle$. b) and c) show in-plane magnitudes (color maps) and polarization textures (polarization ellipses) of $\mathbf{E}_{3k_{\parallel}\sigma\phi_{0}}^{0\parallel}(\mathbf{x}_{\parallel},z=0)$ associated with transitions $|P=3,k_{\parallel}=\textrm{K},\sigma=\pm 1,\phi_{0}\rangle\to|\mathbf{k}^{\parallel}_{f}=\mathbf{0}\rangle$ at $\phi_{0}=0$ and $\phi_{0}=\pi/3$, respectively. Red (blue) polarization ellipses indicate anticlockwise (clockwise) elliptical polarization rotation, and $\mathbf{R}_{0}=-a_{\textrm{BN}}/2$.
It is perhaps natural to consider employing prototypical free electron states that carry well-defined units of intrinsic OAM, e.g., Bessel and Laguerre-Gauss beams [61, 62, 63]. Previous works have demonstrated the ability to monitor OAM exchange between target sample and probing electron through careful pre- and post-selection of the OAM state of the free electrons [51, 42, 52, 43, 44, 53, 45]. However, this strategy is also problematic in the present case, since the reciprocal space wave function densities of these free electron states are azimuthally symmetric, leading to the inability to selectively probe the K or $\textrm{K}^{\prime}$ points. To circumvent this challenge, we introduce pinwheel initial states $|P\,k_{\parallel}\,\sigma\,\phi_{0}\rangle$ of the free electron probe defined as
@@
which are composed of $P$ phased plane wave components with identical in-plane projection $k_{\parallel}=|\mathbf{k}_{\parallel}|$ and phases $\phi_{j}=(2\pi/P)j+\phi_{0}$ winding either counterclockwise ($\sigma=+1$) or clockwise ($\sigma=-1$). The constituent wave vectors for such an incident state are $\mathbf{k}^{i}_{j}=(k^{i}_{\parallel}\cos\phi_{j},\,k^{i}_{\parallel}\sin\phi_{j},\,k^{i}_{z})$, with $k_{i}^{2}=k^{i2}_{z}+k^{i2}_{\parallel}$. For 60 keV probing electrons typical of STEM EEL spectroscopy, the plane waves contributing to the pinwheel superposition states require convergence angles of $\theta_{c}\sim 13.5$ mrad for $k_{\parallel}$ at the hBN K/$\textrm{K}/\textrm{K}^{\prime}$ points, while relative phases may be imprinted using techniques to shape transverse free electron wave functions currently under development [46, 47, 48, 49, 50].
@@
By preparing an incoming pinwheel state and post-selecting the on-axis plane wave state with $\mathbf{k}^{f}_{\parallel}={\bf 0}$, as depicted in Fig. 2a using a spatial filter in the Fourier plane, the measured inelastic scattering process will be composed of $P$ phased momentum recoils with $\mathbf{q}_{j\parallel}=\mathbf{k}^{i}_{j\parallel}$, and the transition field has the form
@@
$\begin{split}\mathbf{E}^{0}_{Pk_{\parallel}\sigma\phi_{0},\mathbf{k}+\mathbf{G}\kappa}=\frac{4\pi ie}{L^{3}}\frac{L}{v_{i}}\frac{\mathbf{k}+\mathbf{G}}{|\mathbf{k}+\mathbf{G}|^{2}}\frac{1}{\sqrt{P}}\sum_{j=1}^{P}e^{\sigma i\phi_{j}}e^{i(\mathbf{k}+\mathbf{G})\cdot(\mathbf{r}_{\kappa}-\mathbf{R}_{0})}\delta_{\mathbf{k}+\mathbf{G},\mathbf{q}_{j}},\\ \end{split}$ (4)

### v2

@@ Abstract
Chirality plays a central role in nearly all aspects of physics, chemistry, and biology. At large length scales, chiral gravitational waves have been considered as sources of circularly polarized contributions to the cosmic microwave background Inomata and Kamionkowski (2019), while at short length scales chirality dictates molecular interactions and organizations underpinning biological processes. In condensed matter physics, chirality and spin-momentum locking are at the heart of the menagerie of identified Hall effects, as well as stable Al Khawaja and Stoof (2001); Yu et al. (2018) and transient free space Gao et al. (2020); Shen et al. (2021) and evanescent Tsesses et al. (2018); Du et al. (2019); Dai et al. (2020); Davis et al. (2020); Ghosh et al. (2023) topological electromagnetic field textures. Following the report Chen et al. (2015) of helicity-resolved Raman scattering experiments from transition-metal dichalcogenide (TMD) atomic layers, it was predicted Zhang and Niu (2015) that two-dimensional (2D) atomic crystals with hexagonal symmetry can host chiral phonons with well-defined pseudo angular momentum (PAM) arising from discrete rotational crystal symmetry. Direct experimental evidence of chiral phonons was subsequently reported based on transient infrared circular dichroism measurements of a monolayer TMD Zhu et al. (2018), generating significant interest in the potential for leveraging chiral phonons to achieve novel material properties that are sensitive to low-energy spin-selective excitations. Beyond 2D hexagonal lattices, chiral phonons have been identified in other 2D Chen et al. (2018, 2019); Suri et al. (2021); Wang et al. (2022); Zhang and Murakami (2022); Fransson (2023) and 3D Chen et al. (2021); Ishito et al. (2023) crystalline systems, as well as in bio-organic molecules Ueda et al. (2023). Chiral phonons have also been recently identified as participants in various exotic condensed matter processes including anomalous thermal Hall conductivity Grissonnanche et al. (2020) and the chiral-phonon-activated spin Seebeck effect Kim et al. (2023), among others Jeong et al. (2022); Chen et al. (2022). Despite the high-level of current interest in chiral phonons, their experimental detection has been largely limited to optical Raman scattering measurements, with spatial resolution and linear momentum limited by the photon wavelength, and direct detection remains an outstanding challenge impeding further progress Wang et al. (2024).
@@
Meanwhile, instrumental advances during the past decade in aberration correction, monochromation, and detector technologies have made it possible to perform vibrational spectroscopy inside a scanning transmission electron microscope (STEM) Krivanek et al. (2014); Dwyer et al. (2016); Lagos et al. (2017); Hage et al. (2018); Hachtel et al. (2019). In particular, momentum-resolved electron energy loss ($q$-EEL) spectroscopy Dwyer et al. (2016); Senga et al. (2019); Hong et al. (2020); O’Hara et al. (2023), and the related high-resolution EEL spectroscopy (HREELS) technique Li et al. (2023, 2024), have been used to measure the energy-momentum dispersion of phonons in mono- and few-layer 2D atomic crystals. While these measurements have demonstrated the ability of inelastic free electron scattering to complement other phonon measurement techniques, such as inelastic neutron and x-ray scattering, the absence of an obvious electron polarization degree of freedom has hindered the spin-selective detection of chiral phonons using free electrons. However, the transverse phase shaping and coherent detection schemes that were originally developed for core-loss EEL spectroscopy and that enabled magnetic circular dichroism measurements in the TEM Yuan and Menon (1997); Hébert and Schattschneider (2003); Schattschneider et al. (2006) have been recently adapted to the low loss regime. By selectively monitoring inelastic scattering of free electrons between states with tailored transverse phase profiles, energy loss processes involving desired target mode symmetries Guzzinati et al. (2017) and OAM states Asenjo-Garcia and García de Abajo (2014); Zanfrognini et al. (2019) have been demonstrated, and optical polarization analogs in free electron scattering have been identified Lourenço-Martins et al. (2021); Bourgeois et al. (2023). The full potential of these measurement schemes is currently being explored, while the transverse phase shaping and selection technologies underlying them are actively under development Niermann et al. (2014)Yu et al. (2023); Feist et al. (2020); Madan et al. (2022); Harvey et al. (2014); Tsesses et al. (2023).
@@
In this Letter we provide a strategy for leveraging inelastic scattering of transversely phase shaped free electrons to achieve direct detection of chiral phonons within a STEM. We introduce pinwheel free electron states composed of a small number of phased plane wave components that share the same reciprocal space three-fold discrete rotational symmetry as chiral phonons in hexagonal monolayer 2D atomic crystals at the K and $\textrm{K}^{\prime}$ points, which introduces a new PAM degree of freedom to the probing electrons. In analogy to previous work on OAM-resolved EEL spectroscopy Lloyd et al. (2012); Asenjo-Garcia and García de Abajo (2014); Cai et al. (2018); Zanfrognini et al. (2019); Lourenço-Martins et al. (2021); Bourgeois et al. (2022, 2023), pre- and post-selection of the transverse free electron states enables tracking of PAM exchange between pinwheel free electrons and the sample. We trace the utility of the identified phase-shaped free electron transitions to the topological spin textures of their associated transition electric fields, which can be tailored to selectively couple to chiral phonons of the same symmetry and PAM. Compared to light-based probes, inelastic electron scattering offers the potential for superior spatial resolution, in addition to an often simpler interpretation since the Brillouin zone (BZ) edges can be accessed directly. This work informs current and future efforts to detect and characterize chiral phonons using inelastic free electron scattering, and identifies specific experimental geometries and parameters to achieve detection.
@@
Figure 1: hBN phonons and $q$-EEL spectroscopy. a) Scheme of $q$-EEL process involving momentum transfer between initial $\mathbf{k}_{i}$ (gray) and final $\mathbf{k}_{f}$ (red) free electron states. The recoil momentum $\hbar{\bf q}$ has projection ${\bf q}_{\parallel}=({\bf q}\cdot\hat{\bf x})\hat{\bf x}$ onto the 2D sample situated in the $xy$ plane. The three $C_{3}$ axes ($C_{3}^{\text{B}}$, $C_{3}^{\text{N}}$, and $C_{3}^{\text{cent}}$) associated with hBN are indicated by red markers in the lower inset. b) Calculated hBN phonon eigenmode dispersion (gray) along $^{\prime}-\Gamma-\textrm{K}$ K and normalized DDCS (color map). c) Real-space in-plane phonon eigenmodes at K and $\textrm{K}^{\prime}$. Pink circles indicate circular motion of sublattices, while arrows show the direction of rotation. The PAM $l^{\text{cent}}_{\lambda}$ is noted below each panel with respect to $C_{3}^{\text{cent}}$. d) Lineouts of the DDCS spectra at K and $\textrm{K}^{\prime}$ from panel (b). All calculations include an empirical damping rate of $\eta=5$ meV/$\textrm{meV}/\hbar$ at zero temperature and a vacuum background.
@@
As a model 2D atomic crystal with hexagonal symmetry that lacks inversion symmetry, we consider chiral phonons in monolayer hexagonal boron nitride (hBN) situated in the $xy$ plane as depicted in Fig. 1a. Eigenenergies $\hbar\omega_{\lambda}(\mathbf{k}_{\parallel})$ and eigenvectors $\bm{\xi}_{\lambda}(\mathbf{k}_{\parallel})$ of the six hBN phonon bands, indexed by $\lambda$, are computed within the harmonic crystal approximation Dove (1993) with force constants taken from the literature Bosak et al. (2007); Michel and Verberck (2008, 2009). The gray lines in Fig. 1b show the eigenenergies as a function of $\mathbf{k}_{\parallel}$ within the first BZ along $^{\prime}-\Gamma-\textrm{K}$ K. In contrast to graphene, the lack of inversion symmetry within the 2D plane and the fact that time reversal symmetry is broken at the K and $\textrm{K}^{\prime}$ points leads to non-degenerate in-plane valley phonon modes at the K and $\textrm{K}^{\prime}$ reciprocal space points with well-defined PAM Zhang and Niu (2015). Real-space motion of the four in-plane phonon eigenmodes ($\lambda=1-4$), at the K and $\textrm{K}^{\prime}$ points are shown in Fig. 1c. Below each image is listed the total PAM $l_{\lambda}^{\text{cent}}$ with respect to the $C_{3}^{\text{cent}}$ rotation axis marked in red at the center of the hexagon shown in the panel (a) inset. Additional details regarding the PAM of in-plane phonons, including definition of PAM relative to other $C_{3}$ axes, is detailed in SM_ . The lowest $(\lambda=1)$ and highest $(\lambda=4)$ energy in-plane modes involve opposing circular motion on both sublattices, with total PAM $l_{\lambda}^{\text{cent}}=0$. For the other two in-plane modes $(\lambda=2,3)$ , while one sublattice remains motionless, the other executes circular motion, the handedness of which ultimately determines whether the PAM $l_{\lambda}^{\text{cent}}=\pm 1$. The depicted phonons exhibit spin-momentum locking in that the $z$-oriented spin associated with each sublattice (equivalent to handedness of circular motion) changes sign under linear momentum reversal, i.e., $\textrm{K}\leftrightarrow\textrm{K}^{\prime}$.
@@
For an incoming probing electron with speed $v_{i}$, the state-resolved loss rate describing inelastic scattering at first order from the 2D atomic crystal at loss energy $\hbar\omega_{if}$ is Hohenester et al. (2018)
@@
where $L^{3}$ defines the probing electron’s quantization volume, $\Omega_{0}$ is the unit cell volume, and $\mathbf{k}$ is within the first BZ. The quantity ~$\tilde{\mathcal{F}}_{\lambda}(\mathbf{k})=\sum_{\mathbf{G}}\sum_{\kappa}\bm{\xi}^{*}_{\lambda\kappa}\cdot M_{\kappa}^{-1/2}\cdot e\bar{\mathbf{Z}}_{\kappa}(\mathbf{k}+\mathbf{G})\cdot\mathbf{E}_{fi,\mathbf{k}+\mathbf{G}\kappa}^{0}$ involves the atomic masses $M_{\kappa}$ and effective charges $eZ_{\kappa}$ of each sublattice, indexed by $\kappa$, and $\mathbf{G}$ denotes a reciprocal lattice vector. $\mathbf{E}_{fi,\mathbf{k}+\mathbf{G}\kappa}^{0}$ is the Fourier coefficient of the quasistatic transition electric field $\mathbf{E}^{0}_{fi}(\mathbf{k})=-i\mathbf{k}(4\pi/k^{2})\,\rho_{fi}(\mathbf{k})$, where $\rho_{fi}(\mathbf{k})$ is the Fourier transform of the charge density $\rho_{fi}(\mathbf{x})=-e\psi^{*}_{f}(\mathbf{x})\psi_{i}(\mathbf{x})$ associated with the free electron transition from initial state $\psi_{i}(\mathbf{x})$ to final state $\psi_{f}(\mathbf{x})$. A phenomenological damping parameter $\eta$ is introduced to give spectral features finite width.
Fig. 1a shows a generic wide field inelastic electron scattering event as typically measured by $q$-EELS whereby an incoming free electron plane wave with wave vector $\mathbf{k}_{i}$ (gray) scatters to an outgoing plane wave with wave vector $\mathbf{k}_{f}$ (red) via interaction with the atomic crystal. In the low loss energy range considered, the linear momentum $\hbar\mathbf{q}=\hbar(\mathbf{k}_{i}-\mathbf{k}_{f})$ lost by the probing electron is transferred into excited phonons with Bloch vector $\mathbf{k}_{\parallel}$. The transition field associated with this situation is SM_
@@
which is linearly polarized along the recoil direction $\hat{\mathbf{q}}$. The rate of phonon mode excitation is dictated by ~$\tilde{\mathcal{F}}_{\lambda}(\mathbf{q})\propto\bm{\xi}^{*}_{\lambda\kappa}\cdot\mathbf{E}^{0}_{\mathbf{q},\mathbf{k}+\mathbf{G}\kappa}\propto\bm{\xi}^{*}_{\lambda\kappa}\cdot\hat{\mathbf{q}}$. Thus, the linearly polarized transition field couples identically to circularly polarized in-plane phonons, and it is not possible to distinguish the PAM of the phonon modes at K and $\textrm{K}^{\prime}$ points from $q$-EEL spectra alone. Indeed, spectral lineouts of the double differential cross section (DDCS) Sakurai (1985); Rossi et al. (2024) at the K and $\textrm{K}^{\prime}$ points from Fig. 1b are shown in Fig. 1d to be indistinguishable.
Figure 2: Transition electric fields involving three-fold pinwheel states. a) Scheme showing a transition from the $|P=3,k_{\parallel},\sigma=+1,\phi_{0}\rangle$ pinwheel state to $|\mathbf{k}^{\parallel}_{f}=\mathbf{0}\rangle$. b) and c) show in-plane magnitudes (color maps) and polarization textures (polarization ellipses) of $\mathbf{E}_{3k_{\parallel}\sigma\phi_{0}}^{0\parallel}(\mathbf{x}_{\parallel},z=0)$ associated with transitions $|P=3,k_{\parallel}=\textrm{K},\sigma=\pm 1,\phi_{0}\rangle\to|\mathbf{k}^{\parallel}_{f}=\mathbf{0}\rangle$ at $\phi_{0}=0$ and $\phi_{0}=\pi/3$, respectively. Red (blue) polarization ellipses indicate anticlockwise (clockwise) elliptical polarization rotation, and $\mathbf{R}_{0}=-a/2\ \hat{\mathbf{y}}$. d) Total PAM $l_{fi}^{i}$ of the transition field relative to the $C_{3}^{i}$ $(i=1,2,3)$ rotation centers marked in panel (b) for $\phi_{0}=0$.
It is perhaps natural to consider employing prototypical free electron states that carry well-defined units of intrinsic OAM, e.g., Bessel and Laguerre-Gauss beams Uchida and Tonomura (2010); Verbeeck et al. (2010); Bliokh et al. (2017). Previous works have demonstrated the ability to monitor OAM exchange between target sample and probing electron through careful pre- and post-selection of the OAM state of the free electrons Lloyd et al. (2012); Asenjo-Garcia and García de Abajo (2014); Cai et al. (2018); Zanfrognini et al. (2019); Lourenço-Martins et al. (2021); Bourgeois et al. (2022, 2023). However, this strategy is also problematic in the present case, since the reciprocal space wave function densities of these free electron states are azimuthally symmetric, leading to the inability to selectively probe the K or $\textrm{K}^{\prime}$ points. To circumvent this challenge, we introduce pinwheel initial states $|P\,k_{\parallel}\,\sigma\,\phi_{0}\rangle$ of the free electron probe defined as
@@
which are composed of $P$ phased plane wave components with identical in-plane projection $k_{\parallel}=|\mathbf{k}_{\parallel}|$ and phases $\phi_{j}=(2\pi/P)j+\phi_{0}$ winding either counterclockwise ($\sigma=+1$) or clockwise ($\sigma=-1$). The constituent wave vectors for such an incident state are $\mathbf{k}^{i}_{j}=(k^{i}_{\parallel}\cos\phi_{j},\,k^{i}_{\parallel}\sin\phi_{j},\,k^{i}_{z})$, with $k_{i}^{2}=k^{i2}_{z}+k^{i2}_{\parallel}$. For 60 keV probing electrons typical of STEM EEL spectroscopy, the plane waves contributing to the pinwheel superposition states require convergence angles of $\theta_{c}\sim 13.5$ mrad for $k_{\parallel}$ at the hBN K/$\textrm{K}/\textrm{K}^{\prime}$ points, while relative phases may be imprinted using techniques to shape transverse free electron wave functions currently under development Niermann et al. (2014)Yu et al. (2023); Feist et al. (2020); Madan et al. (2022); Harvey et al. (2014); Tsesses et al. (2023).
@@
By preparing an incoming pinwheel state and post-selecting the on-axis plane wave state with $\mathbf{k}^{f}_{\parallel}={\bf 0}$, as depicted in Fig. 2a using a spatial filter in the Fourier plane, the measured inelastic scattering process will be composed of $P$ phased momentum recoils with $\mathbf{q}_{j\parallel}=\mathbf{k}^{i}_{j\parallel}$, and the transition field becomes SM_
@@
$\begin{split}\mathbf{E}^{0}_{Pk_{\parallel}\sigma\phi_{0},\mathbf{k}+\mathbf{G}\kappa}=\frac{4\pi ie}{L^{3}}\frac{L}{v_{i}}\frac{1}{\sqrt{P}}\sum_{j=1}^{P}\frac{\mathbf{q}_{j}}{|\mathbf{q}_{j}|^{2}}e^{\sigma i\phi_{j}}e^{i(\mathbf{k}+\mathbf{G})\cdot(\mathbf{r}_{\kappa}-\mathbf{R}_{0})}\delta_{\mathbf{k}+\mathbf{G},\mathbf{q}_{j}}.\\ \end{split}$ (4)